### v2

# Artificial Intelligence AUC Area Under the Curve

Abstract
Initialization plays a critical role in Deep Neural Network training, directly influencing convergence, stability, and generalization. Common approaches such as Glorot and He initializations rely on randomness, which can produce uneven weight distributions across layer connections. In this paper, we introduce the Sinusoidal initialization, a novel deterministic method that employs sinusoidal functions to construct structured weight matrices expressly to improve the spread and balance of weights throughout the network while simultaneously fostering a more uniform, well-conditioned distribution of neuron activation states from the very first forward pass. Because Sinusoidal initialization begins with weights and activations that are already evenly and efficiently utilized, it delivers consistently faster convergence, greater training stability, and higher final accuracy across a wide range of models, including convolutional neural networks, vision transformers, and large language models. On average, our experiments show an increase of $4.9\%$ in final validation accuracy and $20.9\%$ in convergence speed. By replacing randomness with structure, this initialization provides a stronger and more reliable foundation for Deep Learning systems.
1Universitat Politècnica de València
2Universitat Jaume I
3Qsimov Quantum Computing S.L.
11footnotetext: These authors contributed equally.
1 Introduction
Weight initialization remains a critical yet often underappreciated component in the successful training of Deep Neural Networks (including Convolutional Neural Networks and Transformer-based architectures). Although numerous training strategies and architectural innovations receive substantial attention, initialization itself frequently remains in the background, implicitly accepted as a solved or trivial problem. Nonetheless, the initial setting of weights profoundly affects training dynamics, influencing convergence speed, stability, and even the final model’s ability to generalize (LeCun et al., 1998; Glorot and Bengio, 2010; He et al., 2015b).
Historically, weight initialization has evolved significantly. Early methods such as those proposed by LeCun et al. (1998) were primarily heuristic, offering basic variance scaling principles. The seminal work of Glorot and Bengio (2010) rigorously connected weight initialization to variance preservation across layers, introducing a randomized scaling approach that became foundational for Deep Learning (DL). Later, He et al. (2015b) refined this idea, tailoring variance preservation specifically to networks with Rectified Linear Units, and their approach quickly established itself as a default choice in modern DNN training pipelines.
Despite the established efficacy of these stochastic variance-preserving methods, a fundamental question remains largely unchallenged in current literature: Is randomness fundamentally necessary for effective neural network initialization? Random initialization introduces variability that can complicate reproducibility, debugging, and systematic experimentation. Moreover, it implicitly assumes that the optimal starting conditions for training must inherently involve stochasticity, without thoroughly exploring deterministic alternatives that might offer similar or even superior statistical properties.
In this paper, we directly address this unexplored assumption by proposing a novel deterministic initialization scheme, termed Sinusoidal initialization. This technique preserves the critical variance-scaling property central to the effectiveness of Glorot and He initialization schemes, while completely removing randomness from the initialization procedure. It achieves this by employing Sinusoidal functions to deterministically initialize weight matrices, ensuring that each neuron in a given layer receives a distinct weight configuration. This use of Sinusoidal patterns enables rich diversity in weight values while maintaining global variance characteristics, striking a balance between structure and expressive power.
Specifically, our contributions are threefold:
  • We introduce the Sinusoidal initialization, a deterministic initialization strategy utilizing Sinusoidal patterns, explicitly emphasizing symmetry, functional independence, and rigorous variance control for signal propagation, thus eliminating stochastic uncertainties.
  • We present a sound theoretical framework that reveals how stochastic initialization schemes can hinder neuron activation efficiency by inducing asymmetries and imbalances in neuron behavior. By contrast, deterministic approaches that enforce structural symmetry, such as the proposed Sinusoidal initialization, naturally mitigate these issues, leading to more stable activations and improved gradient propagation across layers.
  • Through extensive empirical evaluations involving CNNs, including some efficient architectures, Vision Transformers, and language models, we rigorously demonstrate the empirical superiority of the Sinusoidal initialization over standard stochastic initializations, demonstrating faster convergence and higher final accuracy.
The Sinusoidal initialization consistently yields higher final accuracy and faster convergence, as detailed in Section 4. These benefits are already evident in the case of ResNet-50 trained on CIFAR-100, presented in Figure 1 as a representative example. The experimental setup and results for additional DNN architectures are also provided in Section 4.
Refer to caption Refer to caption
Figure 1: Validation accuracy over training epochs for ResNet-50 on CIFAR-100, comparing Sinusoidal with other initialization schemes.
By challenging the ingrained belief in randomness as a necessary condition for initialization, we open new avenues for rethinking foundational aspects of DNN training, advocating for deterministic and comprehensive practices in the future of DL research.
The remainder of this paper is organized as follows. We begin by revisiting the foundations of weight initialization and its role in DL in Section 2. We then introduce the proposed initialization scheme, outlining its definition and theoretical motivations along Section 3. This is followed by an extensive empirical evaluation in Section 4 across diverse architectures and tasks. Finally, we conclude by summarizing our findings and outlining future research directions in Section 5.
2 Background: Revisiting Initialization
Effective weight initialization is deeply rooted in the principles of variance preservation, which aim to maintain stable signal propagation across layers. This stability is crucial to avoid the notorious problem of vanishing or exploding gradients, which hinder DNN training (Hochreiter et al., 2001).
Formally, for a layer with weight matrix $W\in\mathbb{R}^{m\times n}$, Glorot and Bengio (2010) proposed Glorot initialization, where each weight is sampled from a distribution (either uniform or normal) with zero mean and variance $\textup{Var}(W)=2/(m+n)$, specifically chosen to balance the flow of information during both the forward and backward passes.
Subsequently, Glorot et al. (2011) observed that ReLUs produce naturally sparse activations, typically leaving roughly $50\%$ of hidden units active (positive) and the other 50% inactive (zero). They highlighted the significance of this balanced activation pattern for efficient gradient flow and effective learning. Recognizing this characteristic, He et al. (2015b) refined this initialization specifically for networks using ReLU activations, proposing the variance $\textup{Var}(W)=2/n$ which analytically accounted for ensuring robust data propagation.
While these stochastic methods have been widely successful, their intrinsic randomness introduces significant variability. In particular, the random sampling procedure can sometimes produce weight matrices that lead to suboptimal neuron configurations in the hidden layers. This can cause slow convergence, instability, or even divergence early in training.
Various methods have attempted to reduce or structure this randomness. Orthogonal initialization (Saxe et al., 2014), for instance, ensures orthogonality of weight matrices, stabilizing signal propagation and improving gradient norms, but still involves random orthogonal matrix sampling. Layer-Sequential Unit-Variance (LSUV) (Mishkin and Matas, 2016) takes a step further by adjusting randomly initialized weights iteratively until achieving a precise activation variance. However, LSUV remains data-dependent and partially stochastic, limiting direct reproducibility and practical convenience.
A notable gap thus remains in the literature: Despite these advances, there is currently no widely adopted deterministic initialization method that fully preserves the critical statistical properties of Glorot and He initializations without involving any stochastic component. This unexplored direction motivates our work, wherein we propose a fully deterministic alternative, explicitly designed to overcome these limitations.
3 Definition and Theoretical Foundation of Sinusoidal Initialization
In this section, we introduce a novel deterministic initialization scheme for DNNs, which we refer to as Sinusoidal initialization. This approach is proposed as a theoretically-motivated alternative to conventional stochastic methods, aiming to enable efficient signal propagation and promote functional independence among neurons from the very first training step.
3.1 Sinusoidal Initialization
Let $W\in\mathbb{R}^{m\times n}$ denote the weight matrix of a neural network layer, where $n$ represents the number of input features and $m$ the number of output neurons. We define the weights deterministically using sinusoidal functions as
$W[i,j]=a\cdot\sin(k_{i}\cdot x_{j}+\phi_{i}),$ (1)
where
$k_{i}=2\pi i,\,x_{j}=j/n~{}~{}\textrm{and}~{}~{}\phi_{i}=2\pi i/m,$
for $i\in\{1,2,\dots,m\}$ and $j\in\{1,2,\dots,n\}$.
Under this formulation, each row $W[i,:]$ corresponds to a uniformly sampled sinusoidal waveform of frequency $k_{i}$ and phase offset $\phi_{i}$, evaluated over an evenly spaced grid of $n$ points in the interval $[0,1]$. The linear increase in frequency with the row index ensures that the function $f_{i}(x)=a\cdot\sin(k_{i}\cdot x+\phi_{i})$ completes exactly $i$ full oscillations. The phase offsets $\phi_{i}$ are also chosen to be equally spaced to avoid identical first values across the rows of matrix $W$.
The amplitude $a$ is chosen according to theoretical foundations established by Glorot and Bengio (2010) and He et al. (2015b), which dictate the variance of weights in order to enable efficient signal propagation. Specifically, in our particular case, we adhere to the variance proposed by Glorot, setting the variance of $W$ to be $2/(m+n)$. If $v$ denotes the variance of $W$ when $a=1$, it then follows that $2/(m+n)=a^{2}v$, which yields a natural way to determine the amplitude $a$.
Furthermore, we follow standard practice by initializing all neuron biases to zero, ensuring no preactivation offset is introduced at the start of training.
Weight Distribution
Unlike traditional random initializations, this Sinusoidal scheme yields a non-Gaussian distribution of weights. As illustrated in Figure 3, the weights follow an arcsine distribution, exhibiting higher density near the extrema ±$\pm a$ and lower density near zero. This contrasts with the symmetric bell-shaped curves centered at zero typically produced by standard stochastic initialization methods. Additionally, Figure 3 provides a heatmap of matrix $W$, showcasing the underlying harmonic structure induced by the Sinusoidal construction. The structured and oscillatory pattern of each row is reminiscent of sinusoidal positional encodings employed in Transformer architectures (Vaswani et al., 2017), thereby offering spatial diversity without relying on randomness.
Figure 2: Weight distribution under                                                                                                                                     Sinusoidal initialization.
Figure 3: Heatmap of the                                                                                                                                     initialized weight matrix.
This initialization strategy provides a well-conditioned starting point for early training stages, as will be further discussed in the next subsection.
3.2 Theoretical Motivation
This section provides a theoretical motivation for the design of the proposed initialization. We emphasize its ability to foster efficient training dynamics from the earliest stages. In particular, we analyze the impact that initialization induces on neuron activation patterns in the intermediate layers of DNNs. Furthermore, we introduce a quantitative framework to assess activation quality.
Intuitively, a neuron is inefficient when its activation pattern is imbalanced—i.e., when it remains either active or inactive for a disproportionate amount of samples. Such behavior severely limits its representational power. We formalize this notion through the concept of a skewed neuron:
Definition 1.
Let $Z$ be a random variable representing the output of a neuron with corresponding weights $W_{1},W_{2}\ldots,W_{n}$, and let $\alpha\in(0,1/2)$. The neuron associated with $Z$ is said to be skewed with degree $\alpha$ if
$|P(Z>0\mid W_{1},W_{2},\dots,W_{n})-1/2|>\alpha.$
This definition quantifies the deviation of a neuron’s activation from the ideal $50\%-50\%$ balance. For instance, with $\alpha=0.3$, a neuron is considered skewed if the probability difference between activation and inactivation in absolute value exceeds the $80\%-20\%$ threshold.
A high prevalence of skewed neurons impairs training efficiency by reducing the network’s capacity to exploit nonlinearities and undermining its expressive power, and hence impeding optimization and slowing down convergence.
Empirical evaluation of activation imbalance
To analyze how different initialization schemes affect activation balance, we conduct an experiment with the ViT B16 on the ImageNet-1k dataset. The goal is to compare the extent of activation imbalance across several common initialization strategies, such as Glorot, He, Orthogonal, LSUV, and our proposed Sinusoidal initialization. Figure 4 visually illustrates neuron activation states of the last Feed-Forward of the ViT. Each plot maps neuron indices (horizontal axis) against $768$ randomly selected input samples (vertical axis), with white pixels indicating active (positive output) neurons and black pixels indicating inactive (negative output) ones.
Figure 4: Neuron activation states, in white active neurons ($>0$), for different initializations.
Traditional initializations exhibit distinct vertical white or black bands, indicating neurons that remain consistently active or inactive—i.e., skewed neurons. As noted in Section 2Glorot et al. (2011) emphasized the importance of balanced ReLU activations, with roughly $50\%$ of neurons active to ensure effective learning. Skewed neurons disrupt this balance, limiting the network’s expressive power by underutilizing its nonlinearity and pushing it toward more linear behavior. In contrast, the Sinusoidal initialization yields a more irregular, noise-like pattern with no dominant columns, suggesting a more balanced and expressive activation regime.
To complement this visual analysis, we quantified the proportion of skewed neurons under two thresholds for $\alpha$: $0.1$ and $0.3$. As shown in Table 1, conventional initialization schemes result in an unexpectedly high proportion of skewed neurons, whereas the Sinusoidal approach barely exhibits any skewed neuron. These results highlight the potential of Sinusoidal initialization to better preserve activation diversity. For further experimental details, refer to Appendix C.2.
Table 1: Percentage of skewed neurons under different $\alpha$ thresholds and initialization schemes.
$\alpha$   Glorot He Orthogonal LSUV Sinusoidal
$0.1$ $82.7$ $82.8$ $84.0$ $84.3$ $0.2$
$0.3$ $51.9$ $51.4$ $49.6$ $50.9$ $0.2$
These findings reveal a strikingly high proportion of skewed neurons under conventional random initializations, whereas the proposed deterministic Sinusoidal initialization yields a significantly lower rate. This contrast naturally raises a critical question: What is the reason for this phenomenon? The following subsection aims to shed light on the inner workings of random initializations, why they tend to produce skewed neurons, and why the Sinusoidal scheme successfully mitigates this issue.
Impact of randomness
It is traditionally assumed that randomness in weight initialization is essential for effective training in DNNs. However, the results presented in this work challenge this long-standing belief, revealing that randomness may in fact seed unfavorable configurations from the very beginning.
In a standard random initialization, the weights $W_{1},W_{2},\ldots,W_{n}$ of a neuron are independently sampled from a distribution centered at zero. We identify a critical statistic that governs the initial behavior of the neuron: the cumulative weight imbalance, quantified by the sum
$S=W_{1}+W_{2}+\cdots+W_{n}.$
This quantity captures the imbalance between positive and negative weights. Despite the zero-mean assumption, the value of $S$ can deviate substantially from zero due to statistical fluctuation. For example, under the He initialization (He et al., 2015b), where weights are generated with variance $\sigma^{2}=2/n$, the Central Limit Theorem implies that $S\sim\mathcal{N}(0,2)$.
To empirically investigate the relationship between this statistic and the emergence of skewed neurons, we take, for simplicity, a Multilayer Perceptron (MLP) as the reference model on which to conduct our experiments. For full experimental details, see Appendix C.2. When focusing solely on the first layer of the network, we observe a crystal-clear correlation between the two concepts.
To illustrate this relationship, Figure 5 presents a histogram of $S$ values across all neurons in the first hidden layer. Blue bars correspond to neurons classified as skewed at level $\alpha=0.3$ (i.e. with skewed balance greater than $80\%-20\%$), while orange bars denote non-skewed neurons. The plot reveals a striking pattern: neurons with large absolute values of $S$ are invariably skewed, confirming a strong empirical correlation between $|S|$ and the degree of skewness. This reinforces the central role of $|S|$ in determining activation imbalances at initialization.
Figure 5: Histogram of the statistic $S$ values, highlighting the link between large $|S|$ and neuron skewness at $\alpha=0.3$.
This observation makes it evident that, at the first layer of the model, the relationship between skewed neurons and the corresponding value of the statistic $S$ is remarkably precise: for each value of $\alpha$, there exists a unique threshold $\lambda$ such that a neuron is skewed of degree $\alpha$ if and only if $|S|>\lambda$. This equivalence can be formally established, as stated in the following theorem.
Theorem 1 (Threshold Equivalence).
Let $W_{1},W_{2},\dots,W_{n}$ and $X_{1},X_{2},\dots,X_{n}$ be i.i.d. sequences with
$\mathbb{E}[W_{1}]=0,\quad\mathbb{E}[X_{1}]=\mu>0,\quad\operatorname{Var}(W_{1})=\theta^{2}\in(0,\infty),\quad\operatorname{Var}(X_{1})=\sigma^{2}<\infty.$
Define $S_{n}=\sum_{i=1}^{n}W_{i}$ and $Z_{n}=\sum_{i=1}^{n}W_{i}X_{i}$. Then for any $\alpha\in(0,1/2)$, define
$\lambda_{n}(\alpha):=(\theta\sigma\sqrt{n})\,/\mu\cdot\Phi^{-1}\left(1/2+\alpha\right),$
where $\Phi^{-1}$ is the quantile function of the standard normal distribution. Then, as $n\to\infty$, the following equivalence holds with probability tending to one:
$|P(Z_{n}>0\mid W_{1},W_{2},\dots,W_{n})-1/2|>\alpha\quad\text{ if and only if }\quad|S_{n}|>\lambda_{n}(\alpha).$
This theoretical result tightly links the statistic $S$ with the skewness property of neurons. In the particular case of the Sinusoidal initialization, this connection becomes especially meaningful. Unlike standard random initialization schemes which offer no guarantees on the aggregate behavior of individual neurons, the Sinusoidal initialization exhibits an inherent structural regularity. This symmetry manifests in the exact cancellation of weights along each row of the initialization matrix, ensuring that each neuron begins with a perfectly balanced contribution across its inputs. The following result makes this property precise:
Theorem 2 (Row Cancellation).
Let $W\in\mathbb{R}^{m\times n}$ be defined as in Equation (1). Then, for every row index $i\in\{1,2,\dots,m\}$, the entries along that row sum to zero:
$\sum_{j=1}^{n}W[i,j]=0.$
The proofs and further theoretical analysis can be found in Appendix B.
Consequently, if the objective is to initialize the weights without introducing any biased (i.e., skewed) neurons, then, at least in the first layer, we should aim to enforce $S=0$ for every neuron.
For deeper layers though, the correlation between neuron skewness and the magnitude of $S$ progressively fades due to the asymmetries introduced by preceding layers in the established stochastic initializations. Indeed, the input to an intermediate layer is already biased, and the mean of each input component can deviate more and more, gradually decoupling the notions of skewness and $|S|$. This issue, however, is naturally mitigated by the Sinusoidal initialization. Owing to its strong symmetry, which ensures $S=0$, it guarantees that each neuron begins with an input that is perfectly centered, thereby eliminating any intrinsic directional bias. As a result, the Sinusoidal initialization is theoretically guaranteed to prevent the emergence of skewed neurons, an undesirable behavior frequently observed under stochastic schemes. This conclusion is not only supported by the theorems presented above but also empirically validated in Table 1, where the proportion of skewed neurons is identically zero across all thresholds tested.
The relevance of this symmetry during training will be quantitatively assessed in Section 4. However, before addressing training dynamics, it is crucial to examine a second foundational property: whether individual neurons are initialized with sufficiently diverse and independent functional responses.
Functional independence across neurons
Another critical consideration is the functional independence among neurons in a layer. An optimal initialization scheme should ensure that each neuron focuses on distinct aspects of the input, thereby enhancing representational efficiency.
To detect potential correlations among neurons, we examine horizontal patterns in Figure 4. A dark or white horizontal band would indicate that all neurons respond similarly to a specific input sample, signaling redundancy. As observed, none of the schemes, including the Sinusoidal one, exhibit dominant horizontal patterns. In our case, functional independence is enforced through distinct frequencies $k_{i}$ and phase offsets $\phi_{i}$ for each neuron, evoking the orthogonality of Fourier bases.
In order to further look into more intrinsic horizontal dependencies, the Overfitting–Underfitting Indicator (OUI), as introduced by the authors in  Fernández-Hernández et al. (2025), can be employed to numerically assess functional independence across neurons. This indicator measures the similarity between activation patterns of sample pairs within a given layer, assigning values in the range $[0,1]$, where 0 indicates highly correlated activations and 1 denotes complete independence. In this context, the activation pattern of a sample is understood as a binary vector, where each entry corresponds to the on/off activation state of a hidden neuron. Hence, OUI quantifies how differently pairs of samples activate the network, providing a direct measure of expressive capacity at initialization.
While all methods yield reasonably high OUI scores—Glorot $(0.56)$, He $(0.59)$, Orthogonal $(0.57)$, and LSUV $(0.58)$—the Sinusoidal initialization consistently achieves the highest value, reaching an outstanding $0.98$. This indicates that, although traditional initializations already promote a significant degree of activation diversity, the Sinusoidal scheme provides a strictly better starting point with respect to this metric, offering a more expressive and balanced activation regime from the very beginning.
Theoretical conclusions
The findings of this section lead to two key insights regarding initialization strategies for DNNs:
  1. 1.
    Random initializations naturally induce a high proportion of skewed neurons, which can compromise training efficiency during early stages.
  2. 2.
    Deterministic initializations can substantially mitigate this issue, provided they are designed with a structure that incorporates symmetry, functional independence, and variance control.
From this perspective, the proposed Sinusoidal initialization emerges as a strong candidate: it combines a bounded function family with inherent symmetry, sufficient expressiveness to ensure neuronal independence, and an analytically tractable formulation. Together, these properties make it a promising alternative for establishing a new standard in the initialization of DNNs.
4 Experimental Benchmarking of Sinusoidal initialization
Having introduced the Sinusoidal initialization, a novel deterministic scheme theoretically motivated, we now turn to its empirical validation against both classical and state-of-the-art (SOTA) initialization methods across a range of DNN configurations.
Our experimental validation specifically addresses two main hypotheses: 1) Accuracy improvement: The Sinusoidal initialization achieves higher final validation accuracy compared to alternative initialization schemes in a significant proportion of cases. 2) Accelerated convergence: The Sinusoidal initialization converges faster than any other initialization considered, as measured by the Area Under the Curve (AUC) metric, which besides for classification evaluation, can be used to compare the training convergence speeds (Huang and Ling, 2005).
To evaluate these hypotheses rigorously, we conducted experiments involving five diverse configurations of DNNs and datasets: ResNet-50 (He et al., 2015a) trained on CIFAR-100 Krizhevsky et al. (2009), MobileNetV3 (Howard et al., 2019) trained on Tiny-ImageNet Le and Yang (2015), EfficientNetV2 (Tan and Le, 2021) trained on Tiny-ImageNet, ViT-B16 (Dosovitskiy et al., 2021) trained on ImageNet-1K Russakovsky et al. (2015)), and BERT-mini (Devlin et al., 2019) trained on WikiText Merity et al. (2016). For the first three configurations, we utilized SGD, Adam, and AdamW optimizers, while ViT-B16 and BERT-mini were trained exclusively with SGD and AdamW, respectively, due to the inability of other optimizers to achieve effective training. To ensure a fair comparison between training processes, learning rate schedulers were not used, which may affect the comparability of final accuracies with benchmarks focused on maximizing accuracy.
We trained each configuration fully from scratch using four initialization methods: Default (variants of the He initialization detailed in Table 2), Orthogonal (Orth.), LSUV, and our proposed Sinusoidal initialization (Sin.). The exact training procedures, including hyperparameter settings and implementation details, are provided in Appendix C.3. We measured the maximum validation accuracy achieved at epoch 1, epoch 10, and overall, as well as the AUC for every training scenario. The detailed experimental outcomes are summarized in Table 3.
Table 2: Default initializations for each model.
Model Layer type Default initialization Activation type
ResNet-50 Conv2D He normal ReLU
Linear He uniform
MobileNetV3 Conv2D He normal Hardswish
Linear Normal
EfficientNetV2 Conv2D He normal SiLU
Linear Uniform
ViT-B16 Conv2D Truncated normal ReLU
Linear Glorot uniform GeLU
BERT-mini Linear Normal GeLU
Table 3: Accuracy (%) after 1 epoch, 10 epochs, and at convergence, as well the AUC for each optimizer and initialization scheme (Default, Orthogonal, LSUV, Sinusoidal).
Model / Dataset Optim. 1 epoch accuracy (%) 10 epoch accuracy (%) Maximum accuracy (%) AUC
Def. Orth. LSUV Sin. Def. Orth. LSUV Sin. Def. Orth. LSUV Sin. Def. Orth. LSUV Sin.
ResNet-50 CIFAR-100 SGD $2.6$ $4.0$ $3.6$ $5.0$ $9.7$ $18.2$ $14.8$ $24.3$ $37.3$ $46.5$ $44.8$ $51.9$ $24.7$ $35.1$ $30.9$ $41.8$
Adam $10.1$ $12.5$ $11.8$ $19.9$ $39.9$ $44.1$ $43.1$ $48.7$ $53.1$ $56.6$ $56.3$ $61.5$ $47.9$ $51.2$ $50.9$ $56.5$
AdamW $12.9$ $13.9$ $12.6$ $23.1$ $58.5$ $60.9$ $62.5$ $63.7$ $67.5$ $67.7$ $69.7$ $71.0$ $63.6$ $64.6$ $65.5$ $67.9$
MobileNetV3 TinyImageNet SGD $0.7$ $0.8$ $1.4$ $1.2$ $1.5$ $3.0$ $2.9$ $4.5$ $18.4$ $25.8$ $28.0$ $21.6$ $26.0$ $38.1$ $34.8$ $36.0$
Adam $5.0$ $8.4$ $5.4$ $5.5$ $22.0$ $24.8$ $26.7$ $24.0$ $32.8$ $34.4$ $35.2$ $34.8$ $61.9$ $66.1$ $71.3$ $64.7$
AdamW $13.4$ $14.5$ $14.5$ $12.3$ $40.9$ $43.2$ $40.1$ $42.1$ $40.9$ $43.6$ $40.1$ $42.6$ $79.1$ $84.7$ $76.0$ $82.0$
EfficientNetV2 TinyImageNet SGD $1.0$ $1.2$ $1.9$ $1.0$ $5.0$ $5.3$ $13.4$ $9.0$ $28.1$ $30.9$ $32.1$ $32.0$ $46.7$ $51.9$ $56.4$ $55.6$
Adam $2.9$ $4.2$ $5.6$ $7.2$ $19.3$ $20.2$ $23.0$ $26.4$ $27.7$ $29.8$ $32.7$ $36.6$ $52.9$ $56.1$ $62.4$ $69.5$
AdamW $9.6$ $9.3$ $11.3$ $9.4$ $48.1$ $48.2$ $48.7$ $51.0$ $50.0$ $50.2$ $49.3$ $53.5$ $100.1$ $100.3$ $99.7$ $106.4$
ViT-16 ImageNet-1k SGD $2.8$ $3.8$ $3.5$ $2.6$ $15.2$ $15.2$ $14.1$ $13.9$ $28.6$ $28.2$ $29.6$ $31.5$ $23.3$ $25.2$ $24.1$ $25.4$
BERT-mini WikiText AdamW $9.4$ $6.3$ $11.0$ $6.3$ $13.6$ $10.9$ $14.8$ $17.0$ $40.4$ $42.2$ $15.9$ $41.1$ $58.2$ $71.6$ $32.2$ $72.0$
Figure 6: Training curves showing that Sinusoidal initialization leads to faster convergence compared to default, orthogonal, and LSUV initializations across architectures, using the optimizer that achieved the highest accuracy (SGD for ViT, AdamW for the others).
Analyzing these tables, several clear trends emerge. Firstly, our Sinusoidal initialization consistently achieves higher final validation accuracy compared to other initializations, supporting our first hypothesis. Notably, in the case of MobileNetV3 on TinyImageNet, Orthogonal initialization slightly outperforms our method. Despite this isolated exception, the general trend strongly favors our Sinusoidal initialization. This particular case, however, highlights that this research field remains rich with open avenues for exploration. The behavior observed in this architecture offers a valuable opportunity to further advance our understanding of the DNN training process and refine Sinusoidal initialization strategy accordingly.
Secondly, regarding convergence speed, the Sinusoidal initialization outperforms every other initialization in AUC, except in MobileNet, which indicates a faster convergence and higher accuracy in average over the training process. It also consistently surpasses all other methods in terms of validation accuracy at early stages of training, specifically at epochs 1 and 10, unequivocally confirming our second hypothesis.
Integrating these observations into high-level metrics, our experiments collectively yield an average increase in final validation accuracy of $4.9\%$ over the default initialization, $1.7\%$ over orthogonal and $3.4\%$ over LSUV. Considering the overall training process, the AUC indicates an average improvement of $20.9\%$ over the default initialization, $5.7\%$ over orthogonal and $18.0\%$ over LSUV.
These results strongly support the theoretical predictions from the previous section and underline the practical advantages of the proposed Sinusoidal initialization. Such significant and consistent gains in accuracy and convergence speed emphasize the potential of deterministic, symmetry-driven initialization methods as effective tools for enhancing the performance and efficiency of DNN training.
5 Conclusion
In this article, we introduced and extensively studied the Sinusoidal initialization, a novel deterministic approach to initializing DNNs. Our method fundamentally departs from the traditional paradigm of stochastic initializations, employing structured Sinusoidal patterns designed to enforce symmetry, functional independence, and precise variance control from the outset.
Through rigorous theoretical analysis, we demonstrate that conventional random initialization schemes inherently introduce asymmetries and imbalances, resulting in a high prevalence of what we define as skewed neurons, a notion we formalize and quantify in this work. This skewness substantially undermines neuron activation efficiency and may restrict the expressiveness of the network. In contrast, the deterministic structure of the Sinusoidal initialization is theoretically proven to inherently mitigate these effects.
Empirical evidence strongly supports our theoretical claims. Extensive experiments across diverse DNN architectures, clearly indicate the superiority of Sinusoidal initialization. Specifically, this approach consistently achieved higher final validation accuracy and accelerated convergence rates compared to SOTA stochastic initialization methods. Our results quantify an average accuracy improvement over the default initialization of approximately $4.9\%$ and a notable boost in convergence speed of around $20.9\%$, underscoring the method’s robustness and versatility.
In conclusion, our findings challenge the longstanding assumption that randomness is essential for effective model initialization, presenting strong theoretical and empirical arguments in favor of deterministic methods. The Sinusoidal initialization represents a significant breakthrough in DNN initialization practices, paving the way for future explorations into deterministic initialization schemes, potentially redefining standard practices in the design and training of DNNs.
Acknowledgments and Disclosure of Funding
This research was funded by the projects PID2023-146569NB-C21 and PID2023-146569NB-C22 supported by MICIU/AEI/10.13039/501100011033 and ERDF/UE. Alberto Fernández-Hernández was supported by the predoctoral grant PREP2023-001826 supported by MICIU/AEI/10.13039/501100011033 and ESF+. Jose I. Mestre was supported by the predoctoral grant ACIF/2021/281 of the Generalitat Valenciana. Manuel F. Dolz was supported by the Plan Gen–T grant CIDEXG/2022/013 of the Generalitat Valenciana.
Appendix A Limitations and Societal Impact
This work opens the door to exploring deterministic initializations that exploit symmetry to endow neural networks with desirable properties beyond the reach of stochastic schemes. Among the possible designs, we focused on the Sinusoidal initialization due to its natural structure and analytical tractability. Nevertheless, the broader space of structured deterministic initializations remains largely unexplored, offering a fertile avenue for future research.
In terms of empirical evaluation, while our method consistently outperforms alternatives across most settings, results on MobileNetV3 fall slightly behind those of Orthogonal initialization. Understanding why this architecture deviates from the general trend, despite its apparent similarity to EfficientNetV2, could reveal important insights and guide further refinements. We regard this as a localized irregularity rather than a fundamental shortcoming.
Finally, although we have a clear theoretical understanding of how Sinusoidal initialization guarantees the absence of skewed neurons in MLPs regardless of depth, and unlike traditional stochastic methods, further research is needed to fully characterize this behavior across a broader range of architectures. In particular, establishing a general theoretical framework that ensures the absence of skewed neurons in all established architectures remains an open and promising direction for future work.
Regarding the societal impact, the proposed Sinusoidal initialization enables faster convergence during training, reducing the number of epochs required to reach target performance. This directly translates into lower computational demands, which in turn reduces both the economic cost and the environmental footprint of training deep neural networks.
Given that a substantial portion of the energy consumed in large-scale machine learning pipelines originates from high-performance computing centers, many of which still rely on carbon-emitting energy sources, the adoption of more efficient initialization strategies could contribute to a reduction in overall CO2 emissions. While the exact magnitude of this impact depends on deployment scale and infrastructure specifics, our results highlight a concrete opportunity to align algorithmic efficiency with broader sustainability goals.
Appendix B Theoretical details
This appendix provides the theoretical foundations underlying the results stated in Section 3, and in particular formal proofs of Theorems 1 and 2. We also present a general version of Theorem 1, extending the simplified setting used in the main text to a more flexible framework that covers non-identically distributed inputs and milder moment assumptions.
Beyond these formal results, we include further technical insights that clarify the connection between random initializations and the emergent activation patterns observed in the early layers of deep neural networks.
B.1 Threshold equivalence
In Theorem 1 we stated, under strong i.i.d. assumptions, that the weighted sum $Z_{n}=\sum W_{i}X_{i}$ is controlled with high precision by the magnitude of the total weight $S_{n}=\sum W_{i}$. Specifically, for each level $\alpha\in(0,1/2)$, there exists a sharp threshold $\lambda_{n}(\alpha)$ such that the probability $P(Z_{n}>0\,|\,W_{1},\ldots,W_{n})$ deviates from $1/2$ by more than $\alpha$ if and only if $|S_{n}|>\lambda_{n}(\alpha)$, with high probability.
In this subsection, we show that this result admits a broader formulation under significantly weaker assumptions. In particular, the data $X_{1},\dots,X_{n}$ need not be identically distributed, and only mild regularity is required. The result below extends Theorem 1, and justifies its asymptotic equivalence through general concentration and normalization arguments.
Theorem 3 (Threshold Equivalence, general version).
Let $(W_{1},X_{1}),\dots,(W_{n},X_{n})$ be independent pairs of real random variables, where the weights $W_{1},\dots,W_{n}$ are i.i.d., independent of the data $X_{1},\dots,X_{n}$, and satisfy
$\mathbb{E}[W_{1}]=0,\quad\operatorname{Var}(W_{1})=\theta^{2}\in(0,\infty).$
Assume that the data $X_{1},\dots,X_{n}$ are independent with a common mean $\mu\neq 0$, variances $\operatorname{Var}(X_{i})=\sigma_{i}^{2}\in(0,\infty)$, and that
$\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}^{2}\longrightarrow\bar{\sigma}^{2}\in(0,\infty),\quad\text{as }n\to\infty.$
Finally, assume the conditional Lindeberg condition,
$\frac{1}{\sum_{i=1}^{n}W_{i}^{2}\sigma_{i}^{2}}\sum_{i=1}^{n}\mathbb{E}\left[W_{i}^{2}(X_{i}-\mu)^{2}\mathbf{1}_{\{|X_{i}-\mu|>\varepsilon\sqrt{\sum W_{i}^{2}\sigma_{i}^{2}}\}}\right]\longrightarrow 0\quad\text{for all }\varepsilon>0,$
holds almost surely. Define $S_{n}=\sum_{i=1}^{n}W_{i}$ and $Z_{n}=\sum_{i=1}^{n}W_{i}X_{i}$. Then for any $\alpha\in(0,1/2)$, define
$\lambda_{n}(\alpha):=(\theta\sqrt{\bar{\sigma}^{2}}\sqrt{n})\,/\,\mu\cdot\Phi^{-1}\left(1/2+\alpha\right),$
where $\Phi^{-1}$ is the quantile function of the standard normal distribution. Then, as $n\to\infty$, the following equivalence holds with probability tending to one:
$\left|P(Z_{n}>0\mid W_{1},\dots,W_{n})-1/2\right|>\alpha\quad\text{ if and only if }\quad|S_{n}|>\lambda_{n}(\alpha).$
In particular, for every $\alpha\in(0,1/2)$, the threshold $\lambda_{n}(\alpha)$ is asymptotically unique with high probability.
Sufficient conditions.
The result holds in particular when the data $X_{1},\dots,X_{n}$ are i.i.d. with $\mathbb{E}[X_{1}]=\mu\neq 0$ and $\operatorname{Var}(X_{1})=\sigma^{2}<\infty$. In this case, the conditional Lindeberg condition is automatically satisfied by the classical Lindeberg–Feller Central Limit Theorem (CLT), since the summands $W_{i}X_{i}$ form a triangular array with independent rows and uniformly bounded second moments. These are precisely the assumptions stated in Theorem 1, which is therefore recovered as a particular case of the present general result.
Proof.
Let us write $\Sigma_{n}^{2}:=\sum_{i=1}^{n}W_{i}^{2}\sigma_{i}^{2}$. Given the weights $W_{1},\dots,W_{n}$, we consider the normalized sum
$(Z_{n}-\mu S_{n})/\Sigma_{n}\,\big{|}\,W_{1},\dots,W_{n}.$
Under the assumptions of the theorem, the classical Lindeberg–Feller CLT for triangular arrays implies that this converges in distribution to a standard normal:
$(Z_{n}-\mu S_{n})/\Sigma_{n}\,\big{|}\,W_{1},\dots,W_{n}\xrightarrow{\mathcal{D}}\mathcal{N}(0,1).$
As a consequence,
$P(Z_{n}>0\,|\,W_{1},\dots,W_{n})=\Phi\left(\mu S_{n}/\Sigma_{n}\right)+o(1),$
where the error $o(1)$ vanishes in probability as $n\to\infty$.
Now observe that since $W_{1},\dots,W_{n}$ are i.i.d. with variance $\theta^{2}$, we have by the law of large numbers that
$\sum_{i=1}^{n}W_{i}^{2}=n\theta^{2}(1+o(1)),\quad\text{and}\quad\Sigma_{n}^{2}=\sum_{i=1}^{n}W_{i}^{2}\sigma_{i}^{2}=n\theta^{2}\bar{\sigma}^{2}(1+o(1)).$
Combining these gives
$\frac{\mu}{\Sigma_{n}}=\frac{\mu}{\theta\sqrt{\bar{\sigma}^{2}}}\cdot\frac{1}{\sqrt{n}}(1+o(1)):=c_{n},$
so that
$P(Z_{n}>0\,|\,W_{1},\dots,W_{n})=\Phi(c_{n}S_{n})+o(1).$
Now fix any $\alpha\in(0,1/2)$. Since the standard normal cdf $\Phi$ is strictly increasing and symmetric about zero, we find that
$\left|P(Z_{n}>0\,|\,W_{1},\dots,W_{n})-1/2\right|>\alpha\quad\text{ if and only if }\quad|c_{n}S_{n}|>\Phi^{-1}\left(1/2+\alpha\right).$
Solving for $|S_{n}|$, this is equivalent to
$|S_{n}|>\Phi^{-1}\left(1/2+\alpha\right)/c_{n}=\lambda_{n}(\alpha)(1+o(1)).$
This equivalence holds with probability tending to one, since the approximation error vanishes and $S_{n}$ is independent of the remainder.
Finally, note that the monotonicity of $\Phi$ implies that this inequality can hold for at most one threshold $\lambda_{n}(\alpha)$. Therefore, the threshold is asymptotically unique.∎
Propagation of skewness across layers.
At the input layer, the assumption of equal means across coordinates, central to Theorem 3, is typically satisfied due to standard data normalization practices. However, when standard stochastic initializations are used, the first hidden layer introduces an unavoidable imbalance: the randomness in the initialization causes the activations to deviate from a centered distribution, resulting in the emergence of skewed neurons. Consequently, the inputs to the second layer are no longer mean-centered across dimensions. As training progresses deeper into the DNN, these initial asymmetries are compounded, each layer propagates and potentially amplifies the imbalances introduced by the previous one.
This effect gradually breaks the correspondence between the skewness of a neuron and the statistic $|S|$, as established in Theorem 3. In particular, the conditional expectation $\mathbb{E}[Z_{n}\mid W_{1},\dots,W_{n}]=\mu S_{n}$ hinges critically on the input means $\mu_{i}$ being equal. Once this condition fails, the quantity $\sum_{i}\mu_{i}W_{i}$ governing the conditional bias of the neuron is no longer proportional to $S_{n}$, and the connection between sign bias and the statistic $S$ deteriorates.
This breakdown is clearly observable in the evolution of the histogram shown in Figure 7, which displays the distribution of the skewness statistic $S$ across layers in a multilayer perceptron with four hidden layers. If a layer is initiated with a stochastic criterion (such as Glorot), the distribution of $S$ closely follows a normal distribution, as predicted in Section 3. At the first hidden layer initiated with a random distribution (first row and first column of Figure 7), skewed neurons clearly remain concentrated in the tails. However, as depth increases (second, third and fourth histograms in the first row), these blue tails representing skewed neurons progressively dissolve and spread toward the center of the distribution. As a result, skewed neurons begin to appear even at small values of $S$, indicating that neuron skewness becomes less predictable from the weight imbalance alone.
Figure 7: Evolution of the values of the statistic $S$ for skewed neurons across layers in a MLP.
By contrast, the Sinusoidal initialization fundamentally avoids this pathology. As shown in Theorem 2, each row of the initialization matrix sums to zero, ensuring that the output of the first layer is strictly centered around zero for all neurons (first element of second row). This structural symmetry guarantees that no directional bias is introduced at the very start of the network. More importantly, it prevents the cascade of asymmetries that otherwise accumulates across layers. The rest of the figure shows how Sinusoidal initializations in the first layers of the model help in progressively achieving an equilibrium in terms of skewed neurons. As a result, the relationship between skewness and $S$ remains intact even in deeper layers, and the network preserves maximal discriminatory power throughout its depth. This stability is reflected both theoretically and empirically: in Table 1, the proportion of skewed neurons under Sinusoidal initialization remains identically zero across all tested thresholds. This key structural ingredient, the enforced symmetry at initialization, lays the foundation for a qualitatively new regime in network design. It resolves long-standing inefficiencies related to neuron utilization and enables networks to start from a maximally expressive configuration, thereby accelerating training dynamics systematically. As such, the Sinusoidal initialization marks a turning point in the initialization literature, offering a principled and effective alternative to stochastic schemes.
B.2 Row cancellation in the Sinusoidal initialization
As introduced in Section 3, the Sinusoidal initialization induces a deterministic structure on the weight matrix that stands in stark contrast with the randomness of classical schemes. This structure is not only symmetric, but also highly regular in a way that yields exact cancellation patterns across the rows of the initialization. In particular, each neuron’s incoming weights sum to zero, leading to a strictly balanced preactivation regime from the very first forward pass.
The result we now prove formalizes this cancellation property. To that end, let us recall how the weights are defined under the Sinusoidal initialization. Let $W\in\mathbb{R}^{m\times n}$ be defined coordinate-wise as
$W[i,j]=a\cdot\sin(k_{i}\cdot x_{j}+\phi_{i}),$
where
$k_{i}=2\pi i,\,x_{j}=j/n~{}~{}\textrm{and}~{}~{}\phi_{i}=2\pi i/m,$
for $i\in\{1,\dots,m\}$ and $j\in\{1,\dots,n\}$.
Theorem (Theorem 2 (Row Cancellation)).
Let $W\in\mathbb{R}^{m\times n}$ be defined as in Equation (1). Then, for every row index $i\in\{1,\dots,m\}$, the entries along that row sum to zero:
$\sum_{j=1}^{n}W[i,j]=0.$
Proof.
Fix any $i\in\{1,\dots,m\}$. We consider the sum:
$\sum_{j=1}^{n}W[i,j]=a\sum_{j=1}^{n}\sin\left(2\pi i\cdot\frac{j}{n}+\frac{2\pi i}{m}\right).$
Applying the identity $\sin(\alpha+\beta)=\sin\alpha\cos\beta+\cos\alpha\sin\beta$, we rewrite the sum as
$\sum_{j=1}^{n}\sin\left(2\pi i\cdot\frac{j}{n}+\frac{2\pi i}{m}\right)=\cos\left(\frac{2\pi i}{m}\right)\sum_{j=1}^{n}\sin\left(\frac{2\pi ij}{n}\right)+\sin\left(\frac{2\pi i}{m}\right)\sum_{j=1}^{n}\cos\left(\frac{2\pi ij}{n}\right).$
Thus, to conclude the proof, it suffices to show that both trigonometric sums vanish, that is,
$\sum_{j=1}^{n}\sin\left(\frac{2\pi ij}{n}\right)=0,\quad\sum_{j=1}^{n}\cos\left(\frac{2\pi ij}{n}\right)=0.$
To show this, let $\omega\in\mathbb{C}$ be the unitary complex number and angle $2\pi i/n$. It is clear that $w^{n}=1$, and hence
$\sum_{j=1}^{n}w^{j}=\omega\,\dfrac{1-\omega^{n}}{1-\omega}=0.$
The two real-valued sums above correspond to the imaginary and real parts of this complex sum, respectively. Therefore, both must vanish, and the result follows. ∎
Appendix C Experimental Setup
This appendix contains additional material complementing the main text. Specifically, it provides complete details for all experiments presented in the paper, including training configurations and hardware specifications.
For full reproducibility, all code used to conduct the experiments is publicly available in the accompanying GitHub repository: https://github.com/.
To assess the robustness of our findings, all experiments supporting the main claims were repeated three times. We report mean values across these runs. Standard deviations were computed and found to be consistently small. As these deviations do not materially affect the conclusions, they are omitted from the main tables and figures for clarity.
C.1 Models, datasets and initializations
This work employs a range of DNN architectures across vision and language tasks. The models and datasets used are listed below, along with their sources and brief descriptions.
Models
  • ResNet-50 (Torchvision): A deep residual network with 50 layers, widely used for image classification tasks.
  • MobileNetV3 small (Torchvision): A lightweight convolutional neural network optimized for mobile and low-resource environments.
  • EfficientNetV2 small (Torchvision): A family of convolutional models optimized for both accuracy and efficiency.
  • ViT B16 (Torchvision): A Vision Transformer model that applies transformer architectures to image patches.
  • BERT-mini (Hugging Face): A compact version of BERT for natural language processing tasks, trained with the Masked Language Modeling (MLM) objective.
Datasets
  • CIFAR-100 (Torchvision): A dataset of 60,000 32×32 color images in 100 classes, with 600 images per class.
  • Tiny-ImageNet (Kaggle): A subset of ImageNet with 200 classes and 64×64 resolution images, sourced from the Kaggle repository xiataokang/tinyimagenettorch.
  • ImageNet-1k (Official Repository): A large-scale dataset containing over 1.2 million images across 1,000 categories, used for benchmarking image classification models.
  • WikiText2 (HuggingFace Datasets): A small-scale dataset for language modeling, using the wikitext-2-raw-v1 version to preserve raw formatting.
Initialization Procedures
Four initialization schemes were considered in this work:
  • Default: This corresponds to the initialization procedure provided in the original codebase for each model, as implemented in the respective libraries (e.g., Torchvision or Hugging Face).
  • Orthogonal and Sinusoidal: These initializations are manually applied to the weights of all convolutional and linear layers in the model.
  • LSUV (Layer-Sequential Unit Variance): Implemented using the authors’ original repository (github.com/ducha-aiki/lsuv). This method is applied to convolutional and linear layers and requires a single batch of data from the associated dataset to initialize.
All resources were used in accordance with their respective licenses and usage policies. Note that all dataset splits used correspond to the default configurations
C.2 Activation Imbalance Evaluation
This subsection provides a detailed description of the experiments underlying the theoretical insights discussed in Section 3, particularly those supporting Table 1 and Figures 4 and 5.
Experiment 1: Neuron Activation States (ViT)
To analyze the activation state of neurons, we used the Vision Transformer (ViT-B/16) model trained on ImageNet-1k. The model was initialized using each of the initialization schemes considered in this work, and a forward pass was performed using 768 randomly sampled images from the ImageNet-1k validation set. We captured the output of the second linear layer in the feed-forward block of the last encoder layer—i.e., the final layer before the classifier—resulting in a tensor of shape $768\times 768$ (samples × neurons). To enable visual representation within the paper, we subsampled this tensor to 250 neurons and 250 samples, preserving the structure and interpretability of the activation patterns. This experiment provides the data used in Figure 4.
Experiment 2: Statistic $S$ and skewness Relationship and OUI
To isolate the relation between the statistic $S$ and neuron skewness, we constructed a synthetic feedforward model consisting of a 3-layer MLP with the structure: ReLU → Linear → ReLU → Linear → ReLU → Linear. The input data consisted of random vectors sampled from a Normal distribution. This controlled setting enables clear observation of activation statistics before depth-related effects dilute the signal, as discussed in Appendix B.1.
The histogram in Figure 5 was generated using the outputs from the first linear layer, where the relationship with the statistic $S$ is still prominent. In contrast, the statistics reported in Table 1—such as the percentage of skewned neurons at significance levels $\alpha=0.1$ and $0.3$, and the OUI was computed from the final linear layer in the MLP, where cumulative effects of initialization are more pronounced.
We confirmed the robustness of these results by also sampling inputs from a Uniform distribution, observing consistent trends in skewness behavior and $S$-statistic correlation.
C.3 Benchmark Evaluation
This subsection provides the full experimental details for the training process corresponding to the experiments presented in Section 4, and summarized in Table 3 and Figures 1, 6, and 9. The tables below detail the training hyperparameters (Table 4) and hardware (Table 5) used for each model-dataset pair.
Table 4: Training hyperparameters for all evaluated models.
Model Dataset Epochs Optimizers LR WD Batch size
ResNet-50 CIFAR-100 100 SGD, Adam, AdamW $10^{-3}$ $10^{-3}$ 64
MobileNetV3 Tiny-ImageNet 200 SGD, Adam, AdamW $10^{-3}$ $10^{-3}$ 64
EfficientNetV2 Tiny-ImageNet 200 SGD, Adam, AdamW $10^{-3}$ $10^{-3}$ 64
ViT-B16 ImageNet-1k 100 SGD $10^{-3}$ $10^{-3}$ 64
BERT-mini WikiText 200 AdamW $5\cdot 10^{-5}$ $10^{-3}$ 16
Table 5: Hardware and runtime per single training experiment.
Model Dataset GPU Training time (hours)
ResNet-50 CIFAR-100 NVIDIA A100-SXM4-80GB 2
MobileNetV3 Tiny-ImageNet NVIDIA A100-SXM4-80GB 4
EfficientNetV2 Tiny-ImageNet NVIDIA A100-SXM4-80GB 4
ViT-B16 ImageNet-1k NVIDIA H100-PCIe-94GB 168
BERT-mini WikiText NVIDIA A100-SXM4-80GB 2
Appendix D Additional experimental results
D.1 Arcsine random initialization
To further validate that the key factor behind the effectiveness of our Sinusoidal initialization is the structure it imposes, rather than its unusual distribution, we performed an additional experiment using a randomized initialization. In this variant, weights were sampled independently from the arcsine distribution, which matches the distribution of our Sinusoidal initialization but lacks its balanced structure.
We replicated the experiments from Figure 4 and Figure 5 using this randomized initialization. The results, shown in Figure 8, demonstrate that this initialization behaves similarly to other stochastic methods: the activations exhibit comparable neuron skewness. This confirms that the stochastic nature of the weights, not the distribution, is responsible for the observed effects.
These findings make it clear that the structure introduced by the Sinusoidal initialization is the critical component driving its benefits. This outcome is consistent with the theory established in Subsection 3.2, and further reinforces that the deterministic layout of weights plays a central role in mitigating activation skewness.
Figure 8: Weight distribution, neuron activation states and histogram comparing the statistic S values with neuron skewness at $\alpha=0.3$ for a random initialization with arcsine distribution.
D.2 Additional Training Plots
To complement the main experimental results presented in Section 4, we include in Figure 9 a detailed comparison of the training dynamics between the default and Sinusoidal initialization schemes. For each model and dataset combination, and for every optimizer used during training, we plot both the loss and accuracy curves across epochs.
This visualization allows for a direct assessment of convergence behavior and generalization performance throughout training. Notably, the figure highlights differences in early-stage convergence speed, final validation accuracy levels, and training loss smoothness under the two initialization schemes.
Figure 9: Training loss and accuracy curves for all model-dataset pairs using default and Sinusoidal initialization. Each subplot shows the result for a particular model-optimizer combination.